FTIR AND DFT STUDIES OF NOVEL GERMANIUM-CARBON AND MIXED GERMANIUM-SILICON-CARBON CLUSTERS

\underline{D.L.ROBBINS},C.M.L.RITTBY, and W.R.M.GRAHAM, Department of Physics and Astronomy, TCU, Fort Worth,TX 76129.

The vibrational fundamentals and structures of novel germanium-carbon and mixed germanium-silicon-carbon clusters, formed by laser ablation and trapped in solid Ar, are currently under investigation. ^{a} Vibrational fundamentals and ground state geometries are characterized by comparison of Fourier transform infrared measurements of vibrational frequencies and ^{13}C isotopic shifts for the matrix isolated clusters with the predictions of density functional theory (DFT). Linear GeC_{3}Si has been detected, the first Ge_{j}C_{m}Si_{n} species to be observed optically, and its ν_{1} fundamental identified at 1939.0 cm^{−1}. The results are in excellent agreement with DFT predictions. Additional results and assignments for Ge_{n}C_{m} clusters will be reported.

aD.L.Robbins,C.M.L.Rittby, and W.R.M.Graham J.Chem.Phys. 114, 3570 (2001).